OpenStudy (anonymous):

Consider the equations below for the lengths of line segments AB and AM, where M is the midpoint of line segment AB. AB=9x–3 AM=3x+3 What is the length of line segment AB? AB=

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

AM is half of AB AM = 0.5*AB 3x+3 = 0.5*(9x–3) solve for x